OpenStudy (anonymous):

Are the polygons similar? If they are, write a similarity statement and give a scale factor. The figures are not drawn to scale.

OpenStudy (anonymous):

a. TUWV ~ DEFG; 4.5:2.8 b. TUWV ~ EFGD; 6:2.1 c. TUWV ~ DEFG; 6:4.5 d. The polygons are not similar

OpenStudy (anonymous):

@DarkMoonZ @SyedMohammed98 @undeadknight26 @lerodge857 @sissyedgar @me10000

OpenStudy (anonymous):

@R.Steelheart @chip_lover

OpenStudy (anonymous):

sorry meza no good in math i got a -F sorry :(

OpenStudy (anonymous):

okay :(

OpenStudy (anonymous):

@CausticSyndicalist

OpenStudy (anonymous):

I suck at math. Apologies for failing you D:

OpenStudy (anonymous):

its okay :/

OpenStudy (anonymous):

@josedavid @Legends @serenacox

OpenStudy (josedavid):

yes was up

OpenStudy (anonymous):

the ? at the top i dont know

OpenStudy (legends):

All I know its not D

OpenStudy (anonymous):

:(

OpenStudy (legends):

Lol I don't understand all those other numbers sorry

OpenStudy (anonymous):

its kk :/ and this is only the 2nd ? ctfu

OpenStudy (anonymous):

@sethRB26DETT

OpenStudy (anonymous):

its B hun :D

OpenStudy (anonymous):

OMG THANK YOU SO MUCH CAN I ASK MORE?????????????? PUTTYY PLEEASEEE!!!!!!!!! <3

OpenStudy (anonymous):

wait did you answer yet i was wrong

OpenStudy (anonymous):

No i didnt

OpenStudy (anonymous):

Its c

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

What similarity statement can you write relating the three triangles in the diagram?

OpenStudy (anonymous):

a. RST ~ SRU ~ SUT b. RST ~ RUS ~ SUT c. SUT ~ SRU ~ RTS d. RTS ~ RSU ~ TUS

OpenStudy (anonymous):

What is the value of a and b?

OpenStudy (anonymous):

a = 400/21, b = 580/21 a = 400/21, b = 20/21 a = 580/21, b =29/21 a = 20/21, b = 580/21

OpenStudy (anonymous):

B and A

OpenStudy (anonymous):

ty and it will take a whil to type more but really thank u alot!

OpenStudy (anonymous):

yea no prob hun and dont forget to hit this button-----------^^

OpenStudy (anonymous):

trust me im not

OpenStudy (anonymous):

<3 :)

OpenStudy (anonymous):

What is the value of x to the nearest tenth? The figure is not drawn to scale. a. x=15.7 b. x= 4.4 c. x= 13.9 d. x= 1.9

OpenStudy (anonymous):

6. Are the two triangles similar? How do you know?

OpenStudy (anonymous):

7. Sarah is making a scale drawing of a painting that is 48 in. wide by 120 in. high. Her paper is 12 in. wide and 24 in. tall She decides to use the scale 1 in. = 4 in. Is this a reasonable scale?

OpenStudy (anonymous):

8. Complete the proof. Given: AB dot BE = CB dot BD Prove: ABC~ DBE

OpenStudy (anonymous):

9. If x + 3 / 3 = y + 2 / 3 =??? Explain please.

OpenStudy (anonymous):

10/17 show all steps that you use to solve the problems.

OpenStudy (anonymous):

10. The polygons are similar, but not necessarily drawn to scale. Find the value of x.

OpenStudy (anonymous):

Campsites F and G are on opposite sides of a lake. A survey crew made the measurements shown on the diagram. What is the distance between the two campsites? The diagram is not drawn to scale. The triangles are proportional.

OpenStudy (anonymous):

12. Find the geometric mean of 5 and 8. Write the exact awsner ___ do NOT round.

OpenStudy (anonymous):

13. The measures of the triangle are in the extended ratio 2:2:8. What is the measure of the smallest angle?

OpenStudy (anonymous):

u needed help big time xD

OpenStudy (anonymous):

14. Find the exact length of the altitude drawn th the hypotenuse. Do NOT round. The triangle is not drawn to scale.

OpenStudy (anonymous):

Yes i do:/

OpenStudy (anonymous):

15. What is the solution to the proportion below?

OpenStudy (anonymous):

16. What is the value of x, given that

OpenStudy (anonymous):

17. ABCDE ~ GHJDF. Complete the statement.

OpenStudy (anonymous):

uhhhhhhhhhhhhhh

OpenStudy (anonymous):

im done @sethRB26DETT i promise metal and ive already fanned you! i will do anything just plzz

OpenStudy (anonymous):

lol shut up @chip_lover

OpenStudy (anonymous):

@illkillyou12

OpenStudy (anonymous):

@sleepyhead314

OpenStudy (anonymous):

ITS COOL:)

OpenStudy (anonymous):

@sethRB26DETT MESSAGE ME BACK IN BIO BRUHH